Add
1/4+ 3/10+ 4/15
Divide by moving the decimal point. 27.3 divided by 10 (to the power of 5)


Answer by checkley71(8403)

1/4+3/10+4/15 now find a common denominator.
4=2*2
10=2*5
15=3*5
the 2 & 5 are common to two of the numbers therefore we use them only once & multiply these two 3*3 to the other factors 2*3.
thus we have the LCD of 2*5*2*3=60
(1*15+3*6+4*4)/60
(15+36+16)/60
67/60 answer.
-------------------------------------------------------------------
27.3/10^5
27.3*10^(-5)
.000273 answer.